OpenStudy (anonymous):

Describe 2 methods to convert 9/20 to a decimal

OpenStudy (mertsj):

How do you convert a fraction to a decimal?

OpenStudy (anonymous):

1.you can divide the numerator by the denominator... 9 divided by 20 2.You can get the denominator to equal 100. Then, all you have to do is move the decimal point of the numerator two places to the left, and you have your answer: 9/20 = (5/5)(9/20) = 45/100 Now, move the decimal of the numerator over two places the left to get the final answer: 45/100 = 0.45